OpenStudy (anonymous):

WILL GIVE MEDAL TO BEST ANSWER Mrs. Whimsey's class is having a pizza party. If there are 16 students and 4 pizzas cut into 8 slices, what answer choice represents the fraction of one pizza each student receives? Answers: 1. 1/3 2.1/16 3.1/4 4.1/8

OpenStudy (anonymous):

@Hello_Im_Nick

OpenStudy (anonymous):

Can someone explain that to me? @Hello_Im_Nick @ash2326 @amistre64

OpenStudy (nincompoop):

how about you pay us with food instead?

OpenStudy (anonymous):

@nincompoop Help me?

OpenStudy (anonymous):

actually I think the answer would be 1/4 because 4 pizzas*8 slices=32 and 32/16=2 slices for each student so 2/8 which = 1/4

OpenStudy (anonymous):

Is the question means "each pizza is cut into 8 slices" or "4 pizzas is cut into 8 slices"?

OpenStudy (anonymous):

I think the first is right.

OpenStudy (nincompoop):

sure 4 pizzas cut into 8 that means you would have \[4\times \frac{ 8 }{ 8}\] in total you would have 32 slices if you have 16 students, each student can have 2 slice so we will use 2:16 but that can be simplified to 2/2 : 16/2 1:8

OpenStudy (anonymous):

If this is so, there are 4*8=32 slices of pizza. 32/16=2, each student gets 2 slices. One slice is 1/8 of the pizza, then each student gets 1/4.

OpenStudy (anonymous):

exactly @Chineseboy15

OpenStudy (nincompoop):

if you cut 4 pizzas and yielded only 8 slices then each pizza is only halved if there are 16 students then each student gets a half slice

OpenStudy (nincompoop):

half a slice

OpenStudy (anonymous):

@nincompoop then shouldn't it be 1/16? a pizza

OpenStudy (nincompoop):

yes, mango

OpenStudy (nincompoop):

because the fraction of 1 student gets in a pizza, not of all 4 pizza

OpenStudy (nincompoop):

@Chineseboy15 1/2 is not in the choice but half a slice means half of a half, which amounts to 1/4 but that's a wrong analysis.

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

so the answer is 1/4

OpenStudy (nincompoop):

we are looking how much one would get in A pizza of 8 slices there are 16 students 1 pizza is divided into 8 for a student to get a piece of 1/8, he or she would have to share it with another student so all of them get the same amount

OpenStudy (anonymous):

The answer is 1/4 I think because it is the number of pizzas over the number of student - 4/16 = 1/4

OpenStudy (anonymous):

well @mangorox this is your question you put what you think is right

OpenStudy (nincompoop):